OpenStudy (anonymous):

What is the slope of the line through the points (3, -2) and (-6, -2)?

OpenStudy (anonymous):

Formula Is: y2-y1/x2-x1

OpenStudy (swissgirl):

Ok so the formula is \(\large {y_2-y_1 \over x_2-x_1}\) \((3,-2)=(x_1,y_1)\) \((-6,-2)=(x_2,y_2)\)

OpenStudy (anonymous):

-4/-3

OpenStudy (anonymous):

So: -2-(-2)/-6-3 0/-9 0 So slope is Zero.

OpenStudy (swissgirl):

\(\large { -2-(-2) \over -6-(3)}=\frac{0} {9}\)

OpenStudy (anonymous):

so m=0

OpenStudy (anonymous):

yeah!!!

OpenStudy (swissgirl):

CORRECT :)

OpenStudy (anonymous):

90degree